OpenStudy (anonymous):

State the value of the limit, if it exists lim x→0 sin(x) multiplied with 3x^3+2x^2/x^2 Anybody? :)

OpenStudy (anonymous):

ok this is a big one

OpenStudy (anonymous):

so stay with me

OpenStudy (anonymous):

first we separate the lim

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (3x+2) \times \lim_{x \rightarrow 0}(\frac{ \sin(x) }{ x ^{2} })\]

OpenStudy (anonymous):

i forgot to times x^2 with sin(x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(\frac{ \sin(x)x ^{2} }{ x ^{2} }) \times \lim_{x \rightarrow 0} (3x+2)\]

OpenStudy (anonymous):

ur following the steps right

OpenStudy (anonymous):

wow! :)

OpenStudy (anonymous):

Yep! Think so…

OpenStudy (anonymous):

3x+2 is a polynomial and thus everywhere continuous so \[\lim_{x \rightarrow 0} (3x+2)= 2+3(0)=2\]

OpenStudy (anonymous):

and the \[\lim_{x \rightarrow 0} \frac{ \sin(0) \times 0 }{ 0 } =0\]

OpenStudy (anonymous):

\[ 2 \times 0=0\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

if u have similar denominator like x^2 in ur case u can separate the limit

OpenStudy (anonymous):

Yup! Think so…. :)

OpenStudy (anonymous):

I have to go through it a few times to make sure I understand everything :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Any possibilities of some follow-up questions later on?

OpenStudy (anonymous):

Yea sure

OpenStudy (irishboy123):

you simply can not say this: \[\lim_{x \rightarrow 0}(\frac{ \sin(x)x ^{2} }{ x ^{2} }) \] \[\implies \lim_{x \rightarrow 0} \frac{ \sin(0) \times 0 }{ 0 } =0\] the whole point of this is that dividing by zero you cannot do, so you check what the function in question does as you approach very close to zero. so: \[\lim\limits_{ x→0} \, \, \sin(x) .\frac{3x^3+2x^2}{x^2}\] \[=\lim\limits_{ x→0} \, \, \sin(x) .\lim\limits_{ x→0}\frac{3x^3+2x^2}{x^2}\] \[=\lim\limits_{ x→0} \, \, \sin(x) .\lim\limits_{ x→0}\frac{\frac{3x^3}{x^2}+\frac{2x^2}{x^2}}{\frac{x^2}{x^2}}\] \[=\lim\limits_{ x→0} \, \, \sin(x) .\lim\limits_{ x→0}\frac{3x+2}{1}\] \[=0 \times 2 = 0\]

OpenStudy (anonymous):

Thank you so much for your reply! :)